# Bilevel Reinforcement Learning via the Development of Hyper-gradient without Lower-Level Convexity

### v1

@@ Appendix H Analysis of SoBiRL
$\left\|\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{1}\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{2}\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]\right\|_{2}\leq{HIL_{l}\sqrt{\left|\mathcal{A}\right|}}\left\|\pi^{1}-\pi^{2}\right\|_{2}.$
Denote $z\left(d_{1},d_{2},...,d_{I};x,\pi\right):=l\left(d_{1},d_{2},...,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\widetilde{\nabla}\left(Q_{s_{h}^{i}a_{h}^{i}}-V_{s_{h}^{i}}\right)\left(x,\pi\right)}}\right)$. Taking into account the boundedness and Lipschitz continuity of $l$ in Assumption 6.2, and of ~$\widetilde{\nabla}Q,\ \widetilde{\nabla}V$ in Proposition H.4, we can establish the boundedness,
$\left\|z\left(d_{1},d_{2},...,d_{I};x,\pi\right)\right\|_{2}\leq 2C_{l}HI\frac{C_{rx}}{1-\gamma},$ (70)
and the Lipschitz continuity of $z\left(d_{1},d_{2},...,d_{I};x,\pi\right)$, i.e.,
$\left\|z\left(d_{1},d_{2},...,d_{I};x,\pi_{1}\right)-z\left(d_{1},d_{2},...,d_{I};x,\pi_{2}\right)\right\|_{2}\leq 2C_{l}HI\frac{(1+\gamma)C_{rx}\sqrt{\left|\mathcal{A}\right|}}{\left(1-\gamma\right)^{2}}\left\|\pi_{1}-\pi_{2}\right\|_{2}.$ (71)
@@
~$\widetilde{\nabla}\phi(x,\pi)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]$
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\widetilde{\nabla}\left(Q_{s_{h}^{i}a_{h}^{i}}-V_{s_{h}^{i}}\right)\left({x,\pi}\right)}}\right)\right],$
@@
$\leq$ $\left\|\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{1}\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{2}\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]\right\|_{2}$
$+\tau^{-1}\left\|\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{1}\right)}\left[z\left(d_{1},d_{2},...,d_{I};x,\pi^{1}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{2}\right)}\left[z\left(d_{1},d_{2},...,d_{I};x,\pi^{2}\right)\right]\right\|_{2}$
@@
Under Assumptions 6.2 and 6.3 and given the accuracy in Algorithm 2, we can set the constant step size $\beta<\frac{1}{2L_{\phi}}$, then the iterates $\{x_{k}\}$ satisfy
$\frac{1}{K}\sum_{k=1}^{K}{\left\|\nabla\phi\left(x_{k}\right)\right\|^{2}}\leq\frac{\phi\left(x_{1}\right)-\phi^{*}}{K\left(\frac{\beta}{2}-\beta^{2}L_{\phi}\right)}+\frac{1+2\beta L_{\phi}}{1-2\beta L_{\phi}}L^{2}_{\widetilde{\phi}}\ \epsilon\ ,$
@@
Based on the $L_{\phi}$-Lipschitz smoothness of $\phi(x)$ revealed by Proposition H.6, a gradient descent step results in the decrease in the hyper-objective $\varphi$:

### v2

@@ Appendix H ANALYSIS OF SOBIRL
$\left\|\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{1}\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{2}\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]\right\|_{2}\leq{HIL_{l}\sqrt{\left|\mathcal{A}\right|}}\left\|\pi^{1}-\pi^{2}\right\|_{2}.$
Denote $z\left(d_{1},d_{2},\ldots,d_{I};x,\pi\right):=l\left(d_{1},d_{2},\ldots,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\widetilde{\nabla}\left(Q_{s_{h}^{i}a_{h}^{i}}-V_{s_{h}^{i}}\right)\left(x,\pi\right)}}\right)$. Taking into account the boundedness and Lipschitz continuity of $l$ in Assumption 7.2, and of ~$\widetilde{\nabla}Q,\ \widetilde{\nabla}V$ in Proposition H.4, we can establish the boundedness,
$\left\|z\left(d_{1},d_{2},\ldots,d_{I};x,\pi\right)\right\|_{2}\leq 2C_{l}HI\frac{C_{rx}}{1-\gamma},$ (70)
and the Lipschitz continuity of $z\left(d_{1},d_{2},\ldots,d_{I};x,\pi\right)$, i.e.,
$\left\|z\left(d_{1},d_{2},\ldots,d_{I};x,\pi_{1}\right)-z\left(d_{1},d_{2},\ldots,d_{I};x,\pi_{2}\right)\right\|_{2}\leq 2C_{l}HI\frac{(1+\gamma)C_{rx}\sqrt{\left|\mathcal{A}\right|}}{\left(1-\gamma\right)^{2}}\left\|\pi_{1}-\pi_{2}\right\|_{2}.$ (71)
@@
~$\widetilde{\nabla}\phi(x,\pi)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]$
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\widetilde{\nabla}\left(Q_{s_{h}^{i}a_{h}^{i}}-V_{s_{h}^{i}}\right)\left({x,\pi}\right)}}\right)\right],$
@@
$\leq$ $\left\|\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{1}\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{2}\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]\right\|_{2}$
$+\tau^{-1}\left\|\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{1}\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x,\pi^{1}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{2}\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x,\pi^{2}\right)\right]\right\|_{2}$
@@
Under Assumptions 7.2 and 7.3 and given the accuracy in Algorithm 1, we can set the constant step size $\beta<\frac{1}{2L_{\phi}}$, then the iterates $\{x_{k}\}$ satisfy
$\frac{1}{K}\sum_{k=1}^{K}{\left\|\nabla\phi\left(x_{k}\right)\right\|_{2}^{2}}\leq\frac{\phi\left(x_{1}\right)-\phi^{*}}{K\left(\frac{\beta}{2}-\beta^{2}L_{\phi}\right)}+\frac{1+2\beta L_{\phi}}{1-2\beta L_{\phi}}L^{2}_{\widetilde{\phi}}\ \epsilon\ ,$
@@
[…]